OpenStudy (anonymous):

I need help with this question as soon as possible. Find the vertices of the hyperbola with equation: (x-1)^2 / 4 - (y+2)^2 / 9 = 1

OpenStudy (anonymous):

(x,y)=(1,-2)

OpenStudy (anonymous):

Thank you :)

OpenStudy (amistre64):

the equation gives you all the needed clues. -(-1,2) is the center: (1,-2) the "x" part is positive so it rides along parallel to the x axis; meaning we adjust the "x" part from the center and 4 is the under side of our "x" part so sqrt it to find the distance between vertexes" sqrt(4) = 2 from center we go 1 to the left and 1 to the right (1+1 = 2) (1+1,-2) and (1-1,-2) (2,-2) (0,-2) let me wolfram to dbl chk it tho

OpenStudy (amistre64):

i was a bit off in my remembering; sqrt(4) = 2 then adjust left and right from center by "2" ....

OpenStudy (amistre64):

(1+2,-2) and (1-2,-2) (3,-2) (-1,-2) we could make it simpler by lifting the equation to the x axis .. but thats for another time i spose